PartnerObserving Log for 2007 June 09/10

Technical Observer: J. Hill
Telescope Operator: D. Gonzalez
Instrument: LBC-Blue

Technical observing with a bit of SDT M81 science at the beginning of the night.

Opened at sunset. T=11.9degC. Wind SW at 3 m/sec D=-7.6degC. Very clear, although smoky below the boundary layer at ~12000 feet. Thin clouds in the very far West.

5 bias

02:57 SkyFlatR_5 x4

03:03 SKyFlatRrot180_5 x10 (32000,.....

Editted phu_sci.hdr and corresponding tech files to correct comment about TELAZ convention in headers.

03:08 Move to field of M81_extra2 for focussing (but, waiting for darkness....)

PG1633_UB (have not been focussed yet, so 10 arcsec donuts)

PG1633_Vgr

03:30 Back to M81, el=43deg

(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
03:33 Vfastextra 3937 0 0 0 0 0 sent

03:35 Vfastextra -141 -164 -619 -532 106 -27 sent

03:37 Vfastextra -65 -29 -150 -147 20 35 sent

03:39 Vfastextra -72 -262 -35 -17 -42 41 sent

03:42 Vfinefoc - likes -0.048 at 3.72

03:46 Vcofenif - likes -0.080 at 3.76 (cofenif is finefoc reversed) Does this mean there is hysteresis?

03:53 M81BV - guidestar is astigmatic, seeing was ~1 arcsec while focussing, but worse than 2 arcsec in data.

04:06 Vfastextra -1468 146 24 39 124 -65 sent

04:08 M81V - Images 1.1 arcsec at 1.55 airmasses.

Move to M101_NE

04:19 Vfastextra -722 -268 272 321 -404 -44 sent

04:21 Vfastextra -87 -155 -171 53 -40 103 sent

04:24 Vfinefoc - likes 0.00 at 4.31

04:28 Vcofenif - likes -0.002 at 3.37 So no apparent hysteresis here.

04:33 Bfinefoc - likes 0.08 at 5.58 (ratty curve)

04:37 Vfinefoc - likes -0.011 at 4.31

04:43 Vfastextra -371 219 462 49 35 118 sent

IIF::init: CSQ cannnot be found among subsystems that respond to 'trigger' message [N] IIF::IIF(const char, const char) : IIF::init() failed; please consult earlier syslog messages for more details Error on TCSInitialize()

04:47 Vfastextra -136 151 -203 134 -47 -24 sent

04:50 start of M101_Ctr, stopped after first two dithers.

Move to 16:00 +44:00

05:02 Vfastextra -184 -736 355 -31 -106 -51 sent (was this astigmatism noise from a bad pupil?)

05:04 Vfastextra 151 -632 -134 10 29 53 sent

Move to 16:00 +44:10

05:07 Vfastextra -176 675 188 -52 -23 -78 sent

05:09 Vfastextra 25 31 290 -6 129 -46 sent

05:13 Rfastextra -29 390 105 78 -134 138 not sent

05:15 Rfinefoc - likes +0.081 at 3.16

05:21 Gfinefoc (had a mirror failed to focus because of wrong state) - liked 0.086 but don't trust it.

05:25 Bfinefoc - likes +0.043 at 3.86 (all the defocussed images are astigmatic -- like field dependent astigmatism that LBCFPIA is not fully correcting at the field center)

05:29 Vfastextra -193 242 -420 140 -70 79 sent

05:32 Vfastextra -170 -337 -67 -19 41 -77 sent

05:36 Gfinefoc - likes +0.0838 at 3.04

05:43 Rfinefoc - likes +0.088 at 2.80

05:47 Gfinefoc - likes 0.0836 at 3.05

05:52 Rfinefoc (mirror fails to focus) - liked 0.094 at 3.03 but don't trust it

05:58 Vfastextra -176 450 -93 125 -239 -11 sent

06:02 Rfinefoc - likes +0.0967 at 3.30

06:06 Bfinefoc - likes +0.0673 at 3.20

06:11 Sfinefoc - likes +0.440 at 3.67 (but squirrelly curve)

06:15 Rfinefoc - likes +0.0672 at 2.89 (drifted?)

06:19 Vfastextra -443 315 208 -76 155 -8 sent

06:22 Vfastextra 61 150 -86 31 -25 -56 sent

06:24 Rfinefoc - likes +0.1016 at 2.68

06:29 Sfinefoc - likes +0.4588 at 3.34

06:33 Ufinefoc - likes +0.3527 at 3.39

06:39 Vfinefoc - likes +0.0327 at 3.17

06:43 Vfastextra 151 207 -1 156 41 -6 sent

06:46 Ufinefoc - likes +0.356 at 4.29

06:50 Sfinefoc - likes +0.4549 at 3.48

06:54 Bfinefoc (mirror failed to focus at posn 3) - likes +0.0524 at 3.69 (suspect)

Move to field of GRB070518, at EL=67.5, using new r-SLOAN focus offset of +0.088

07:01 Vfastextra -596 -940 172 -129 -42 -46 (note huge astigmatism correction)

07:03 Vfastextra 27 414 -71 93 -13 27

07:06 GRB070518 - first dither is 3.0 pix (0.68 arcsec) at center and 3.4 pix at chip 4 at 1.08 airmasses.

Note that rotator angle is -180 for this case. Seems to confirm that the field-dependent astigmatism is from the hub or below and not from the cryostat.

I noticed tonight (and previously) that the "Stellar FWHM" reported on the LBC UI from the guide stars seems to reach a floor of ~1.3 arcsec. While the science images are 0.7 to 0.8 arcsec and also the tech images are near 0.8. See lbcbtec.20070610.071641_1 (and many images nearby) for an example. (sent email to Gallozzi)

07:27 GRB070518 - second pass

07:48 Vfastextra -614 -460 85 149 70 -6 sent

RXg=+40 Yg=1.784 (Why were the global offsets previously zero?)

1270,3188 moves to 1351,3341 with this adjustment at ROTATOR angle -200 (a shift of 39 arcsec)

07:53 Vfastextra -131 66 -449 380 54 26 sent

07:56 GRB070518 - third pass (with different collimation and thus different pointing)

Had a soft UI hang during dither 3.

For this dither 1 image at rotator angle -200 the images are worse at the bottom of the array. So probably my collimation correction was too big.

08:19 RXg=+20 Yg=+0.892 (but note that Ya=-0.418 cancelling some of the Y correction)

08:20 Vfastextra -184 328 -28 -83 3 48 sent

08:23 GRB070518 - fourth pass (with half the previous collimation adjustment)

I can't identify any certain gradient in the up/down image quality at rotator angle -220. Maybe chip 1 is 0.1 pixel worse? Images are 0.6 arcsec FWHM in the center and 0.7 arcsec FWHM at the edge.

20:38:31 +21:12:04

08:57 B1sec on pointing star EL=64, Rot=-58 star at 1071,2989 [center is 1048.5, 2957.8]

17:46:27.51 +27:43:14

09:05 B1sec on pointing star EL=73, Rot=68 star at 968,2957

19:06:14.9 -04:52:57

09:14 B1 sec on pointing stat EL=52, Rot=+1 star at 1018,3016

Edit LBCLEFTpoinitngmodel.dat

IE from -41.3 to -54.9 stopped and restarted PCS on lbtmu101

09:23 B1sec star at 1016,2965

IE to -56.86 IA from -279.62 to -272.85 stopped and restarted PCS on lbtmu101

09:28 B1sec star at 998,2957 sign of IA correction wrong

IA to -286.40 stopped and restarted PCS on lbtmu101

09:32 B1sec star at 1033,2962 EL=52 Rot=+7

IA to -289.70 IE to -57.86

09:36 B1sec star at 1039,2963

CA from +24 to +4 arcsec/radian

09:40 B1sec star at 1125, 2979 (move mostly toward +X on the detector)

CA back to its normal +24

20:45:17.3 +61:50:19

09:55 B1sec star at 1004, 2971 EL=59, Rot=-156

Move to open cluster at 20:34 +28:18, EL=83

10:08 Vfastextra -790 97 1197 585 222 -95 sent

10:11 Vfastextra 370 380 131 -17 -51 -46 sent

10:12 Vfastextra -52 370 -111 -38 57 3 sent

10:16 Rfinefoc - likes +0.0824 at 2.91, but images have a load of astigmatism/trailing

10:22 Bexexinin

10:27 Clear C00

10:29 Vfastextra -740 268 226 1056 -9 360 sent

10:32 Vfastextra -249 -290 -257 -794 114 416 sent

10:36 Vfastextra -295 241 -166 -276 -29 41 sent

10:39 Rfinefoc - likes +0.0855 at 2.6, images are still not very pretty out of focus, astigmatism at the bottom of chip 2. Seeing must be less than 0.5 arcsec, since these images are dominated by the astigmatism.

Ya=-0.234 Yt=0.778 Yg=+0.892 But which is the sign of astigmatism that we need atEL=82?

10:44 Vexexinin

10:59 Bexexinin (mirror failed to focus)

Clear the Y, RX values from global offsets

11:09 Vfastextra -97 -457 -295 -28 -182 -119 sent

11:11 Vfastextra -31 -152 70 -29 -41 -85 sent

11:14 Rfinefoc (mirror failed to focus on image 2?) likes 0.0819 at 2.58 with ellipticity ~0.2, Needs more study to say if we moved the global offset in the correct direction or not.

11:23 Vallextra (telescope moved just before it finished - observer error)

11:30 Close dome T=8.9degC, D=-7.5degC Wind calm, Sky clear.

5 bias

Focus Offsets

Adopted Filter Focus Offsets relative to measurements with Vfastextra using alpha=76.64.

V-BESSEL == adjusted to be 0.000

B-BESSEL returned to near its historical value = +0.053

U-BESSEL adjusted to = +0.353

r-SLOAN increased to +0.086

g-SLOAN = 0.083

SDT_Uspec refined to = 0.454

These offsets have been installed in leftchannelfocus.dat, but not yet in the various other places that they need to go.

Comparison of old and new versions of LBCFPIA

From last night using image lbcb.20070610.111126.fits
11:11 Vfastextra -31 -152 70 -29 -41 -85 - median diameter 61.14 pixels

From the LBCFPIA version of lbcfpia_20061105
11:11 Vfastextra 162 -152 70 -29 -41 -85 - median diameter of 61.14 pixels

So this is not the cause of the apparent focus shift between ~November 2006 and ~March 2007. The measured diameter is the same, and the shift in measured Z4 is the difference in calibration of "alpha".

-- JohnHill - 10 Jun 2007
Topic revision: r8 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback